A person with myopia has a far point of 2.4 m. This person wears glasses to correct their vision.

If the person wears their glasses 1.5 cm from their eye, what is the optical power of the corrective lenses that allow them to see distant objects clearly? (to 2 s.f and in D)
1
Expert's answer
2020-05-25T11:16:21-0400
"P=\frac{1}{d_0}+\frac{1}{d_i}"

"P=\frac{1}{\infty}+\frac{1}{0.015-2.4}=-0.42\ D"